OpenStudy (anonymous):

Solve 2x – 5 = 3(2x + 5)?

OpenStudy (anonymous):

2x-5=3(2x+5) 2x-5=6x+15 (distributed the 3) 2x=6x+20 (added the 5) -4x=20 (subtracted the 6x) x=-5

OpenStudy (anonymous):

then u can put the -5 back in the equation to make sure the two sides equal one another

OpenStudy (anonymous):

thanks man and ok